What is the natural wet bulb globe temperature to which a construction worker, working outside in the sun, is exposed if the following temperatures are obtained: Dry bulb = 90 °F; Natural wet bulb = 86 °F; Globe = 96°F? One of equations you may use: WBGT = 0.7 WB + 0.2 GT + 0.1 DB (with solar load); WBGT = 0.7 WB + 0.3 GT (without solar load); WBGTaverage = ? (WBGTi * Ti) / ? (Ti).
Answer #1

WBGT – Wet Bulb Globe Temperature

WB – Wet Bulb Temperature

GT – Globe Temperature

DB – Dry Bulb Temperature

under conditions of direct sunlight (outdoors, no cloud cover)

WBGT = 0.7 WB + 0.2 GT + 0.1 DB

            = 0.7*86 + 0.2*96 + 0.1*90

            = 60.2 + 19.2 + 9

            = 88.4 F

in environments that are indoors, in the shade, or on a cloudy day

WBGT = 0.7 WB + 0.3 GT

            = 0.7*86 + 0.3*96

            = 60.2 + 28.8

            = 89 F

For different temperatures and respective WBGTs are substituted in the expression to find average value.